lkjhgfd:

Mandy obtains a $155,000 20/6 balloon mortgage with a rate of 4.25%. What will her monthly payments be? (2 points)

justus:

The formula for fixed monthly mortgage repayment formula is A= P * r * \[\frac{ r(1+r)^n }{ (1+r)^n-1 }\]

justus:

A is the payment amount per month, R is the interest rate is 4.25 %, N is number of months which is 240 months, P is mortgage value which is $155,000.

justus:

A= (0.0425)(155000) \[\frac{ 0.0425(1+0.0425)^240 }{ (1+0.0425)^240-2 }\]

justus:

This will now be $279.98.

justus:

Hope this helps you out.